OpenStudy (anonymous):

A stone is dropped from the edge of a cliff. What is the velocity of the stone in m/s after 3 seconds?

OpenStudy (anonymous):

Just asking, do you know what is velocity?

OpenStudy (anonymous):

velocity is distance over time right?

OpenStudy (anonymous):

well that is speed, but yeah

OpenStudy (anonymous):

good enough

OpenStudy (anonymous):

yes that is right, thinking about physics

OpenStudy (anonymous):

how do i know what the distance is?

OpenStudy (anonymous):

what force is acting on the stone?

OpenStudy (anonymous):

gravity?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

so how does gravity tie in with the formula

OpenStudy (anonymous):

the acceleration of gravity can be approximated at -9.8m/s^2

OpenStudy (anonymous):

so it gravity the distance? i'm still confused

OpenStudy (anonymous):

you can multiply the velocity by the seconds. 9.81(3)=29.43ms y=5(9.81)(3)carot sign(2), g is gravity (9.81m/s carot sign2) y=5(9.81)(3) carot sign 2 y=44.145 It is just the same as the last one and just use 4 instead of 3 7=78.48m

OpenStudy (anonymous):

so the velocity is 78.48?